# The trade-off between diagonal and off-diagonal elements in the eigenstate thermalization hypothesis

### v2

@@ Appendix A Variance relation
Here, we present the detailed derivation of the key quantum variance relation given in eqs. 13 and III.3 of the main text. Starting with the definition
$V(\rho_{B},O^{B_{1}}_{ij})=\lVert(\sigma^{ij}_{B_{1}}-\delta_{i,j}\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$=\text{Tr}(\sigma^{ij}_{B_{1}}\rho^{-1}_{B_{1}}\sigma^{ij{\dagger}}_{B_{1}})-\delta_{i,j}.$ (34)
Assuming the spectral decomposition $\rho_{B_{1}}=\sum_{\alpha}p^{\prime}_{\alpha}\Pi^{\alpha}_{B_{1}}$, we can express
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\text{Tr}(\sigma^{ij}_{B_{1}}\Pi^{\alpha\beta}_{B_{1}})\rvert^{2}-1,$ (35)
where $\Pi^{\alpha\beta}=\ket{\alpha}\bra{\beta}$ |. Recognizing that $\lvert\text{Tr}(\sigma^{ij}_{B_{1}}A^{B_{1}})\rvert^{2}=\lvert\text{Tr}(\Pi^{ij}A^{B_{1}})\rvert^{2}$, and using
$\mathcal{J}_{\rho_{B_{k}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{k}})=(p^{\prime}_{\alpha}p^{\prime}_{\beta})^{-1/2}\Pi^{\alpha\beta}_{B_{k}}$ (36)
@@
$O^{B_{1}}_{\alpha\beta}=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{1}}),$ (37)
we can further rewrite
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\text{Tr}(\Pi^{\alpha\beta}_{B_{1}}\Pi^{ij}_{B})\rvert^{2}-1$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{1}}_{\alpha\beta})^{\dagger}\Pi_{i}O^{B_{1}}_{\alpha\beta}]-1$
$=\sum_{\beta}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\rho_{B_{1}}O^{B_{1}}_{ii})-1=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (38)
which directly corresponds to the relation in eq. 13.
The proof of section III.3 follows similar approach as appendix A. The key difference is the substitution of $O$ with $\overline{O}$, and the use of the relation
$\lvert\text{Tr}[(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}})A^{B_{1}}]\rvert^{2}=\lvert\text{Tr}[\Pi^{ij}(\frac{1}{\mathcal{C}}\sum_{k}A^{B_{k}})]\rvert^{2}.$ (39)
Here, we need to use $\overline{O}^{B}_{\alpha\beta}=\frac{1}{\mathcal{C}}\sum_{k}O^{B_{k}}_{\alpha\beta,1}$, where $O^{B_{k}}_{\alpha\beta,1}$ represents the translational copies of $O^{B_{1}}_{\alpha\beta}$. Using this, we can rewrite the expression as
$V(\rho,\overline{O}^{B_{1}}_{ii})+\sum_{j;j\neq i}V(\rho,\overline{O}^{B_{1}}_{ij})$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(\overline{O}^{B}_{\alpha\beta})^{\dagger}\Pi_{i}\overline{O}^{B}_{\alpha\beta}]-1$
$=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,k}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{k}}_{\alpha\beta,1}]-\frac{1}{\mathcal{C}}$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{l}}_{\alpha\beta,1}]-\frac{\mathcal{C}-1}{\mathcal{C}}.$ (40)
@@
$\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{k}}_{\alpha\beta,1}]$
$=\sum_{\beta}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\rho_{B_{1}}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ii,k}_{B_{1}}))=\text{Tr}(\rho_{B_{1}}^{-1}).$ (41)
@@
Figure 6: Parameters: $g=1.05$, $N=15$. (a) The number of states of the spin chain system. (b) Energy probability distribution of the canonical ensemble with $\beta=0.1$. (c) Energy as a function of the thermodynamic beta. (d) The value of $\log_{10}I(B_{1},B_{2})$ between two sites at $\beta=0.1$.
Appendix B Basic properties of the model
@@
Here, we briefly illustrate the basic properties of the model. As shown in fig. 6(a), the density of states for both integrable and non-integrable systems is quite similar, with the non-integrable system exhibiting a smoother distribution that can be reasonably approximated by a binomial distribution [14]. The energy-temperature curve for the system is provided in fig. 6(c). Additionally, as shown in fig. 6(d), the mutual information $I(B_{1},B_{2})$ decays exponentially with the distance between sites. The plot depicts the system with $h=0.1$; the mutual information for the $h=0$ system is nearly identical and therefore not shown.

### v3

@@ Appendix A Variance relation
Here, we present the detailed derivation of the key quantum variance relation given in Eqs. 22 and III.3 of the main text. Starting with the spectral decomposition of the reduced density matrix $\rho_{B_{1}}$ within its Hilbert space framework, this operator admits the standard representation
$\rho_{B_{1}}=\sum_{\alpha}p^{\prime}_{\alpha}\Pi^{\alpha}_{B_{1}},$ (44)
where $\{\Pi^{\alpha}_{B_{1}}\}$ constitutes a complete set of orthogonal projection operators. This spectral structure allows us to reformulate Section II.1 as
$V_{\text{dg}}^{i}+\sum_{j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\alpha}}^{-1}\lvert\bra{\beta}\sigma^{ij}_{B_{1}}\ket{\alpha}\rvert^{2}-1$
$=\sum_{j,\alpha,\beta}{p^{\prime}_{\alpha}}^{-1}\lvert\text{Tr}(\sigma^{ij}_{B_{1}}\Pi^{\alpha\beta}_{B_{1}})\rvert^{2}-1,$ (45)
where the elementary operators $|\Pi^{\alpha\beta}_{B_{1}}=\ket{\alpha}_{B_{1}}\!\bra{\beta}$ | span the operator space of the subsystem, forming a complete basis for operator expansion. Recognizing that $\lvert\text{Tr}(\sigma^{ij}_{B_{1}}A^{B_{1}})\rvert^{2}=\lvert\bra{E_{j}}A^{B_{1}}\ket{E_{i}}\rvert^{2}$, we can further rewrite
$V_{\text{dg}}^{i}+\sum_{j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\alpha}}^{-1}\lvert\bra{E_{j}}\Pi^{\alpha\beta}_{B_{1}}\ket{E_{i}}\rvert^{2}-1$
$=\sum_{\alpha,\beta}{p^{\prime}_{\alpha}}^{-1}\bra{E_{i}}\Pi^{\beta\alpha}_{B_{1}}\Pi^{\alpha\beta}_{B_{1}}\ket{E_{i}}-1$
$=\sum_{\alpha}{p^{\prime}_{\alpha}}^{-1}\bra{E_{i}}I_{B_{1}}\ket{E_{i}}-1=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (46)
where in the second equality we use the orthogonal projection completeness condition $\sum_{j}\ket{E_{j}}\bra{E_{j}}=I$. The Appendix A directly corresponds to the relation in Eq. 22. Another derivation method that is easy to generalize to average observable needs to use the following relationship from Eq. 44
$\mathcal{J}_{\rho_{B_{k}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{k}})=(p^{\prime}_{\alpha}p^{\prime}_{\beta})^{-1/2}\Pi^{\alpha\beta}_{B_{k}},$ (47)
@@
$O^{B_{1}}_{\alpha\beta}=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{1}}).$ (48)
Using these two equations and $O^{B_{1}}_{\beta\alpha}=(O^{B_{1}}_{\alpha\beta})^{\dagger}$, we can prove that
$V_{\text{dg}}^{i}+\sum_{j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\alpha}}^{-1}\lvert\bra{E_{j}}\Pi^{\alpha\beta}_{B_{1}}\ket{E_{i}}\rvert^{2}-1$
$=\sum_{\beta,\alpha}p^{\prime}_{\beta}\text{Tr}[O^{B_{1}}_{\alpha\beta}\Pi^{i}_{B}(O^{B_{1}}_{\alpha\beta})^{\dagger}]-1$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{1}}_{\alpha\beta})^{\dagger}\Pi^{i}_{B}O^{B_{1}}_{\alpha\beta}]-1$
$=\sum_{\beta,\alpha}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\Pi^{\beta\alpha}_{B_{1}}\Pi^{i}_{B}\Pi^{\alpha\beta}_{B_{1}})-1=\text{Tr}(\rho_{B_{1}}^{-1})-1.$ (49)
The proof of Section III.3 follows similar approach as Appendix A. Different from the $V(\rho_{B},O^{B_{1}}_{ij})$ considered in Appendix A, what we need to consider here is $V(\rho_{B},\overline{O}^{B_{1}}_{ij})$ with $\overline{O}^{B_{1}}_{ij}=\frac{1}{\mathcal{C}}\sum_{k}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ij,k}_{B_{1}})$. The key difference is the substitution of $O$ with $\overline{O}$. Following Appendix A and using the relation
$\lvert\text{Tr}[(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}})\Pi^{\alpha\beta}_{B_{1}}]\rvert^{2}=\lvert\text{Tr}[\Pi^{ij}(\frac{1}{\mathcal{C}}\sum_{k}\Pi^{\alpha\beta,1}_{B_{k}})]\rvert^{2}$ (50)
with $\Pi^{\alpha\beta,1}_{B_{k}}$ represents the translational copy of $\Pi^{\alpha\beta}_{B_{1}}$, we can reach a formula similar in Appendix A
$V(\rho_{B},\overline{O}^{B_{1}}_{ii})+\sum_{j\neq i}V(\rho_{B},\overline{O}^{B_{1}}_{ij})$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(\overline{O}^{B}_{\alpha\beta})^{\dagger}\Pi^{i}_{B}\overline{O}^{B}_{\alpha\beta}]-1,$ (51)
where $\overline{O}^{B}_{\alpha\beta}=\frac{1}{\mathcal{C}}\sum_{k}O^{B_{k}}_{\alpha\beta,1}$ and $O^{B_{k}}_{\alpha\beta,1}=\mathcal{J}_{\rho^{1}_{B_{k}}}^{-1/2}(\Pi^{\alpha\beta,1}_{B_{k}})$ represents the translational copies of $O^{B_{1}}_{\alpha\beta}$. Go on with Appendix A, we have
$V(\rho_{B},\overline{O}^{B_{1}}_{ii})+\sum_{j\neq i}V(\rho_{B},\overline{O}^{B_{1}}_{ij})$
$=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,k}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi^{i}_{B}O^{B_{k}}_{\alpha\beta,1}]-\frac{1}{\mathcal{C}}$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k\neq l}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi^{i}_{B}O^{B_{l}}_{\alpha\beta,1}]-\frac{\mathcal{C}-1}{\mathcal{C}}.$ (52)
@@
$\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi^{i}_{B}O^{B_{k}}_{\alpha\beta,1}]$
$=\sum_{\beta}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\Pi^{\beta\alpha}_{B_{k}}\Pi^{i}_{B}\Pi^{\alpha\beta}_{B_{k}})=\text{Tr}(\rho_{B_{1}}^{-1}).$ (53)
@@
Figure 3: Parameters: $g=1.05$, $N=15$. (a-b) The number of states of in a spin chain under PBC (OBC). (c) Energy probability distribution of the canonical ensemble with $\beta=0.1$. (d) Energy as a function of the thermodynamic beta. (e-f) The value of $\log_{10}I(S_{i},S_{j})$ between two sites $S_{i}$ and $S_{j}$ at $\beta=0.1$ in a spin chain under PBC (OBC).
Appendix B Numerical Verification of ETH
B.1 Basic properties of the model
@@
Here, we briefly illustrate the basic properties of the model. As shown in Fig. 3(a), the density of states for both integrable and non-integrable systems is quite similar, with the non-integrable system exhibiting a smoother distribution that can be reasonably approximated by a binomial distribution [14]. The energy-temperature curve for the system is provided in Fig. 3(c). Additionally, as shown in Fig. 3(d), the mutual information $I(B_{1},B_{2})$ decays exponentially with the distance between sites. The plot depicts the system with $h=0.1$; the mutual information for the $h=0$ system is nearly identical and therefore not shown. From the perspective of these basic properties, integrability has no obvious impact on them.
B.2 The impact of energy difference and temperature on distinguishability
Figure 4: Parameters: $g=1.05$, $N=15$, $N_{B_{1}}=1$. (a) Local diagonal element measure $V_{\text{dg}}^{i}$ as a function of energy deviation $\mathcal{E}:=E_{i}-E(\rho_{B})$, where $\rho_{B}$ is fixed as a maximally mixed state $\rho_{B}=\sum_{i}\Pi^{i}_{B}/D_{N}$ with $E(\rho_{B})=0$. For each situation, only a subset of 600600600 states with approximately uniform energy spacing is plotted, rather than all states across the energy range. (b) The average distinguishability $\overline{V}^{\text{Loc}}_{\text{dg}}$ within the energy shell $\mathcal{M}_{E,\Delta=0.2}$. For each energy shell centered at $E$, the $\rho_{B}$ in $V_{\text{dg}}^{i}=V(\rho_{B},O^{B_{1}}_{ii})$ is adaptively parametrized as $\rho^{\text{c}}(\beta)$ to satisfy $\text{Tr}[H\rho^{\text{c}}(\beta)]=E$.
Although equation (22) holds for all energy eigenstates, the measure $V_{\text{dg}}^{i}$ for the diagonal elements varies across different energy eigenstates |$\ket{E_{i}}$. This difference is evident in Fig. 1. In more detail, as shown in Fig. 4(a), there is some randomness in the measure $V_{\text{dg}}^{i}$ with respect to the eigenstate, but generally, the greater the energy difference between $E_{i}$ and $E(\rho_{B})$, the higher the average distinguishability. In addition, integrability also has a great influence. For the integrable systems, the diagonal elements $V_{\text{dg}}^{i}$ tend to be more variable, and the average distinguishability is typically higher.
$\mathcal{M}_{E,\Delta}=(E-\Delta/2,E+\Delta/2).$ (55)
Moreover, when analyzing energy eigenstates within $\mathcal{M}_{E,\Delta}$, it is preferable to use an ensemble $\rho_{B}$ with the same energy as $E$. In Fig. 4(b), we selected the canonical ensemble with corresponding energies for different energy shells and calculated the average distinguishability as
$\overline{V}^{\text{Loc}}_{\text{dg}}=\sum_{E_{i}\in\mathcal{M}_{E,\Delta}}V_{\text{dg}}^{i}/d_{\mathcal{M}_{E,\Delta}},$ (56)
The results in Fig. 4 are consistent with the traditional ETH: As the energy discrepancy between the ensemble and a particular energy eigenstate grows, their distinguishability increases correspondingly. Furthermore, the distinguishability can only be minimized when the ensemble is specifically tailored to match the energy of a given eigenstate.
B.3 The ETH under the local measurement
In this section, we focus on analyzing the strong and weak ETH under local measurements. We examine two energy shells $\mathcal{M}^{\prime}:=\mathcal{M}^{\prime}_{E=0,\Delta}$ and $\mathcal{M}:=\mathcal{M}_{E(\beta),\Delta}$, where $\Delta=0.2N$ and $E(\beta)=E(\rho^{\text{c}}(\beta=0.1))$. Both energy shells include highly excited states. For the energy shell $\mathcal{M}$, we considered the canonical ensemble $\rho^{\text{c}}(\beta=0.1)$ and the mirocanonical ensemble $\rho^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}}\Pi^{i}/d_{\mathcal{M}}$. For the energy shell $\mathcal{M}^{\prime}$, we used the mirocanonical ensemble ${\rho^{\prime}}^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}^{\prime}}\Pi^{i}/d_{\mathcal{M}^{\prime}}$.
$\overline{V}^{\text{Loc}}_{\text{(off-)dg}}=\frac{1}{d_{\mathcal{M}_{E,\Delta}}}\sum_{E_{i}\in\mathcal{M}_{E,\Delta}}V_{\text{(off-)dg}}^{i}.$ (57)
$V^{\text{Loc}}_{\text{dg},\max}=\max_{E_{i}\in{\mathcal{M}_{E,\Delta}}}V_{\text{dg}}^{i}.$ (58)
$V^{\text{Loc}}_{\text{off},\max}=\max_{E_{i}\in{\mathcal{M}_{E,\Delta}};j}V_{\text{off}}^{ij}.$ (59)
Here, as in Eq. 22, the index $j$ spans the full Hilbert space.
Figure 5: The local quantity $\text{Tr}(\rho_{B_{1}}^{-1})-1$ under different system sizes. Parameters: $g=1.05$, $N_{B_{1}}=1$. In the figure, zero markerstrokewidth corresponds to PBC, whereas non-zero markerstrokewidth indicates OBC.
According to Eqs. 22 and 32, the local state of ensemble $\rho_{B_{1}}$ is a crucial quantity. As $N$ increases from 101010 to 151515, according to Fig. 5, $\text{Tr}(\rho_{B_{1}}^{-1})-1$ shows almost no variation across several ensembles, regardless of whether the system is integrable.
Figure 6: Parameters: $g=1.05$, $N_{B_{1}}=1$. In the figure, zero markerstrokewidth corresponds to PBC, whereas non-zero markerstrokewidth indicates OBC. Linear fits are shown exclusively for the energy shell $\mathcal{M}_{E(\beta),\Delta}$ with $\rho_{B}=\rho^{\text{mc}}$. The exponential decay constant is estimated using this linear fit. (a) The average diagonal elements $\overline{V}^{\text{Loc}}_{\text{dg}}$ versus $N$. Exponential decay constants: Under PBC, $0.76\times\ln 2$ (standard error $0.05\times\ln 2$) for the non-integrable system and $0.22\times\ln 2$ (standard error $0.05\times\ln 2$) for the integrable system. Under OBC, $0.84\times\ln 2$ (standard error $0.03\times\ln 2$) for the non-integrable system and $0.12\times\ln 2$ (standard error $0.02\times\ln 2$) for the integrable system. (b) The maximum diagonal element $V^{\text{Loc}}_{\text{dg},\max}$ versus $N$. Exponential decay constants: Under PBC, $0.51\times\ln 2$ (standard error $0.13\times\ln 2$) for the non-integrable system and $0.00\times\ln 2$ (standard error $0.03\times\ln 2$) for the integrable system. Under OBC, $0.30\times\ln 2$ (standard error $0.03\times\ln 2$) for the non-integrable system and $-0.06\times\ln 2$ (standard error $0.04\times\ln 2$) for the integrable system. (c) The average off-diagonal elements $\overline{V}^{\text{Loc}}_{\text{off}}$ versus $N$. Both non-integrable and integrable systems exhibit similar exponential decay constants of $1.00\times\ln 2$ with standard errors $<0.001\times\ln 2$. (d) The maximum off-diagonal element $V^{\text{Loc}}_{\text{off},\max}$ versus $N$. Exponential decay constants: Under PBC, $0.34\times\ln 2$ (standard error $0.14\times\ln 2$) for the non-integrable system and $0.17\times\ln 2$ (standard error $0.08\times\ln 2$) for the integrable system. Under OBC, $0.78\times\ln 2$ (standard error $0.09\times\ln 2$) for the non-integrable system and $0.11\times\ln 2$ (standard error $0.04\times\ln 2$) for the integrable system.
To avoid cluttering the figures, we perform linear fitting only for the microcanonical ensemble $\rho^{\text{mc}}$ with energy shell $\mathcal{M}_{E(\beta),\Delta}$; all subsequent discussions of exponential decay constants are based on this case. As shown in Fig. 6(a-b), there is a significant difference between integrable and non-integrable systems in the behavior of the diagonal elements. In the non-integrable system, the average distinguishability of the diagonal elements decreases exponentially with $N$. The exponential decay constant is approximately $0.76\times\ln 2$ for PBC and $0.84\times\ln 2$ for OBC, which is closely aligns with the $\ln 2=d\ln(1/D_{N})/dN$ derived in Eq. 23. The exponential decay constant for the maximum $V_{\text{dg}}^{i}$ is slightly smaller, around $0.51\times\ln 2$ for PBC and $0.30\times\ln 2$ for OBC. In contrast, for the integrable system, the exponential decay constant for the average distinguishability is minimal, about $0.22\times\ln 2$ for PBC and $0.12\times\ln 2$ for OBC. The maximum $V_{\text{dg}}^{i}$ seems to have almost no decay with $N$ for both boundary conditions. As shown in Fig. 6(c-d), the difference in off-diagonal elements between integrable and non-integrable systems is small. Consistent with the prediction of Eq. 23, the average off-diagonal elements are strictly suppressed exponentially, with an exponential decay constant of $1.00\times\ln 2$ in both cases, regardless of integrability. However, the exponential decay in the average off-diagonal elements does not imply that the maximum values of these elements will also decrease exponentially. As shown in Fig. 6(d), although the maximum off-diagonal values do decrease with $N$ from the fit: Under PBC, about $0.34\times\ln 2$ for the non-integrable system and about $0.17\times\ln 2$ for the integrable system, they do not exhibit a clear pattern. Under PBC, there is a clear distinction between the integrable and non-integrable cases, and the exponential decay behavior is more pronounced. The exponential decay constant is about $0.78\times\ln 2$ for the non-integrable system and about $0.11\times\ln 2$ for the integrable system.
Additionally, the average off-diagonal elements of the integrable system in Fig. 6(c) are always slightly smaller than those of the corresponding non-integrable system. This observation is consistent with the fact that the average diagonal elements of the integrable system in Fig. 6(a) are always larger than the average diagonal elements of the corresponding non-integrable system and the trade-off relationship between the diagonal and off-diagonal elements described in Eq. 32.
Moreover, through comparative analysis of PBC and OBC, we observe distinct thermalization characteristics in non-integrable systems. Under OBC configurations, the maximum values of off-diagonal elements in the ETH framework exhibit pronounced exponential decay behavior, a phenomenon likely attributable to the breaking of energy degeneracy. Furthermore, both the average and maximum diagonal elements of ETH demonstrate universally larger magnitudes in OBC systems compared to their PBC counterparts. This can be explained by the absence of the boundary interaction term $-\sigma_{z}^{1}\otimes\sigma_{z}^{N}$, which impedes efficient information propagation through the system. The resulting suppression of information diffusion creates additional barriers to thermalization, making the equilibration process more challenging in OBC configurations compared to systems with periodic boundaries.
B.4 The ETH under the average observable
Figure 7: Parameters: $g=1.05$, $N_{B_{1}}=1$. Linear fits are shown exclusively for the energy shell $\mathcal{M}_{E(\beta),\Delta}$ with $\rho_{B}=\rho^{\text{mc}}$. The exponential decay constant is estimated using this linear fit. (a) The average diagonal elements $\overline{V}^{\text{Avg}}_{\text{dg}}$ versus $N$. Exponential decay constant are $0.71\times\ln 2$ (standard error $0.05\times\ln 2$) for the non-integrable system and $0.18\times\ln 2$ (standard error $0.03\times\ln 2$) for the integrable system. (b) The maximum diagonal element $V^{\text{Avg}}_{\text{dg},\max}$ versus $N$. Exponential decay constants are $0.35\times\ln 2$ (standard error $0.07\times\ln 2$) for the non-integrable system and $0.19\times\ln 2$ (standard error $0.03\times\ln 2$) for the integrable system. (c) The average off-diagonal elements $\overline{V}^{\text{Avg}}_{\text{off}}$ versus $N$. Both non-integrable and integrable systems exhibit similar exponential decay constants of $1.1\times\ln 2$ with standard errors $<0.01\times\ln 2$. (d) The maximum diagonal element $V^{\text{Avg}}_{\text{off},\max}$ versus $N$. The non-integrable system shows a decay constant of $0.31\times\ln 2$ (standard error $0.26\times\ln 2$). For the integrable case, the decay rate is statistically insignificant (fitted value non-positive due to strong fluctuations) and thus omitted.
Now, let us explore the strong and weak ETH under the average observable. The energy shells and ensembles are consistent with those in the previous section. Given that the ensembles used here are all translation-invariant, we simply replace $V(\rho_{B},O^{B_{1}}_{ij})$ with $V(\rho_{B},\overline{O}^{B_{1}}_{ij})$ to derive the corresponding $\overline{V}^{\text{Avg}}_{\text{(off-)dg}}$ and $V^{\text{Avg}}_{\text{(off-)dg},\max}$ from Eqs. 57, 58 and 59. According to Sections III.3 and 37, the local term is obtained by multiplying $\text{Tr}(\rho_{B_{1}}^{-1})-1$ from the previous section by a factor of $1/\mathcal{C}$. In the case of the average observable, the spatial correlation of the ensemble states, represented by $V^{\text{Avg}}_{\text{Cor}}$, also needs to be considered. As shown in Fig. 8, the correlation term is negative, which further lowers the upper bound of the average (off-)diagonal elements. Additionally, the correlation terms vary significantly depending on the energy center, as the ensemble $\rho^{\text{mc}}$ with energy near zero exhibits weaker correlations. This difference also explains why the two ensembles in Fig. 7(c) are more different from those in Fig. 6(c): although the local states of the two ensembles are very similar, their correlation terms are different, which will cause the difference between the two ensembles in Fig. 7(c) according to Eq. 37. However, according to Eq. 32, the difference in correlation terms cannot be reflected, resulting in almost no difference between the two ensembles in Fig. 6(c).
As shown in Fig. 7(a-d), the significant difference in the exponential decay constants of the diagonal elements between integrable and non-integrable systems remains evident. Due to the additional attenuation introduced by the factor $1/\mathcal{C}$ in Eq. 37, the average of the off-diagonal elements is exponentially suppressed, with an exponential decay constant even larger than $\ln 2$. The maximum values of the off-diagonal elements still show no clear pattern. However, in Fig. 7(c), the average off-diagonal elements of the non-integrable system are no longer consistently smaller than those of the corresponding integrable system. This observation remains consistent with the trade-off relationship between the diagonal and off-diagonal elements given by Eq. 37, as Fig. 7(a) shows that the average diagonal elements of the integrable system are now longer always larger than those of the non-integrable system.